# supplementary

### v1

@@ Acknowledgments
JB has received research funding from Bayer AG and travel expenses from Merck KG outside of the presented work. The funders had no role in study design, data collection and analysis, publication decisions, or manuscript preparation. ODP received personal honoraria from Merck Sharp & Dohme and Almirall, received travel support from Kyowa Kirin, Pierre Fabre, Sanofi and Sun Pharma and is member of the advisory board of Bristol Myers Squibb and Sanofi. These entities had no role in study design, data collection and analysis, publication decisions, or manuscript preparation. All other authors declare no conflicts of interests.
@@
LS, CL, JB, KB designed the study. CL, DH, JB provided samples and clinical data for NucSeg, TumSeg and the not-curated H&E nucleus dataset. DH, ODP, SD, AK, ML, AF, JL provided samples and clinical data for the CPI dataset. LS, CL, DH, JB generated and cured annotations for NucSeg and TumSeg datasets. LS conceived Histo-Miner software as well as its evaluation. CL and JB conceived the immunotherapy response case study. LS conducted statistical analyses and methods for immunotherapy response case study.
@@
{adjustwidth}
30pt30pt where $I$ input Image containing $N$ pixels, $p_{i}(I;\mathbf{w_{0},w_{1}})$ regression output of HoVer branch at pixel $i$, $w_{0}$ and $w_{1}$ 2 sets of weights. $\Gamma_{i}$ defines the groundtruth of the horizontal and vertical distances of nuclear pixels to their corresponding centers of mass, the horizontal and vertical components of this map are denoted $\Gamma_{i,x}$ and $\Gamma_{i,y}$ respectively (see [14] for visualization and definitions). m denotes total number of nuclear pixels within the image and M denotes the set containing all nuclear pixels. $\nabla_{x}$ and $\nabla_{y}$ denote the gradient in the horizontal x and vertical y directions respectively. Finally, $X$ denotes the branch groundtruth, $Y$ the branch prediction, K is the number of classes and $\epsilon$ is a smoothness constant that was set to $10^{-3}$.
@@
On Table.1 the different hyperparameters tested are displayed. $cat-max-ratio$ corresponds to the max area ratio that could be occupied by single category for a given crop of the input image. $img-scale$ corresponds to the resizing of the image before cropping. These resizing can randomly be modified by a factor contained in [0.5, 2], following the data augmentation pipeline from the semantic segmentation library MMSegmentation [42].
Not all combinations were fully tested, some were aborted if the training loss did not decrease in the firsts epochs or if several hyperparameters in the set were already poorly performing in previous trainings.
@@
Suppl. Fig 1: Example of distance overestimation with search area being a square. Here we have 1 cell in the squared search area $S$ and 2 cells in $(C\setminus S)$. While calculating the minimum distance of cells from the center $O$, if we only consider cells in $S$, we will take $cellA$ as being the closest cell. In reality the closest cell is $cellB$. The distance calculated will be the distance between $cellA$ and the center, and will then be higher than the real closest distance. Nevertheless, the overestimation is bounded as expressed in Eq. 22 with $a=1$ as we are here in the case of a squared search area.
@@
Suppl. Fig 2: Probability of having distance overestimation representations, with search area being a square. (a) During the first search, if the search box is a square of side lenght 2r, we could have a distance calculation error if a cell is in the blue square $S_{1}$ without being in the pink circle $C_{1}$ (then in $S_{1}\setminus C_{1}$), and another cell is in black circle $C_{2}$ without being in in the blue square $S_{1}$ (then in $C_{2}\setminus S_{1}$). Then the cell in $C_{2}\setminus S_{1}$ could be missed from closest distance calculation and the cell in $S_{1}\setminus C_{1}$ considered as the closest cell. In other cases, such as having at least one cell in $S_{1}$, no distance overestimation could be made. (b) To be more precise, any cell in the hatched blue area inside $S_{1}\setminus C_{1}$ monitored by $x$ would lead for sure to an error if another cell is in the hatched black area inside $C_{2}\setminus S_{1}$ monitored by $x$ For now we focused on the probability of overestimation in the case that at least one cell is found during the first search. If no cell is found, then the calculation differs for the further searching steps as visualized in (c) & (d). In such cases, a specific new area cannot contain any cells, as none were found during prior searches, impacting the full calculation.
@@
$B_{N}:(\text{There is among the cells in $(C_{2}\setminus S_{1})$ a cell closer to the center $O$ than all cells in $S_{1}$})$
@@
$\bar{B_{N}}:(\text{All the cells in $(C_{2}\setminus S_{1})$ are further away to the center $O$ than all cells in $S_{1}$})$
@@
$V_{N}(n,d):(\text{There are $N-n$ cells in $S_{1}$ and the closest cell to theorigin $O$ is at a distance $d$ from $O$})$
@@
$\tilde{V}_{N}(n,d):(\text{There are $N-n$ cells in $S_{1}$ and they are all located at a distance greater than $d$ from $O$})$
@@
$\operatorname{area}(hatchblack)$ is the area of 4 circular segments of arc radius $d$ and sagitta ($d$ $r_{1}$) then:
@@
{adjustwidth}
30pt30pt where $P(A_{N})$ is calculated in (11), $P\bigl{(}U_{N}(0)\bigr{)}$ is calculated in (10), $P\bigl{(}U_{N}(n)\bigr{)}$ is calculated in (9) and finally $P\bigl{(}\bar{B_{N}}\mid U_{N}(n)\bigr{)}$ is developed in (13), (14), and (20).
@@
{adjustwidth}
30pt30pt where $w$ is the width of the rectangle and $a$ the aspect ratio.
We covered here the case where at least one cell is found in the first search, meaning at least one cell is in $S_{1}$. In practice, distances can be overestimated in subsequent search iterations following the first. Unfortunately the calculation will be slightly different. Visualization of the cases where no cell is found inside $S_{1}$ are available at Supplementary Fig. 2(c), 2(d).
@@
In addition to figure Fig.4.b form the core test, we provide visualization of ROC curve for each splits.

### v2

@@ Acknowledgments
JB has received research funding from Bayer AG and travel expenses from Merck KG outside of the presented work. The funders had no role in study design, data collection and analysis, publication decisions, or manuscript preparation. ODP received personal honoraria from Merck Sharp & Dohme and Almirall, received travel support from Kyowa Kirin, Pierre Fabre, Sanofi and Sun Pharma outside of the presented work and is member of the advisory board of Bristol Myers Squibb and Sanofi. These entities had no role in study design, data collection and analysis, publication decisions, or manuscript preparation. All other authors declare no conflicts of interests.
Ethics approval and consent to participate
The study was performed in agreement with the Declaration of Helsinki Institutional Review Board of the University Hospital Bonn (vote number 187/16), Ethics committee of the University Hospital Cologne (votenumbers 21–1500, 20–1082 and 22–1330-retro) and institutional review board of the TU Munich (vote number 2024–363-S-CB - 1). Need for informed consent was waived for this retrospective analysis using anonymized data.
@@
LS, CL, JB, KB designed the study. CL, DH, JB provided samples and clinical data for NucSeg, TumSeg and the not-curated H&E nucleus dataset. DH, ODP, SD, AK, ML, AF, JL provided samples and clinical data for the CPI dataset. LS, CL, DH, JB generated and cured annotations for NucSeg and TumSeg datasets. LS conceived Histo-Miner software as well as its evaluation. CL and JB conceived the immunotherapy response case study. LS conducted statistical analyses and methods for immunotherapy response case study. LS, JB, KB reviewed the manuscript.
@@
Poor precision in tumor and healthy epithelial detection from state of the art pretrained models
Experts and learning algorithm have hard time discriminating between tumor epithelial and healthy epithelial in the context of cSCC skin cancer, especially when not enough context is provided. Both cell classes have very close morphologies, and need a broader context to be separated. We used Hovernet [14], CellViT-256 and CellVit-SAM-H [30] models pretrained on Pannuke dataset [15] to apply inference on 2 TumSeg samples where tumor regions are annotated by 2 experts. We show in Supplementary Fig. 1 that the models are not able to recognize healthy epithelial, and classify most of the epithelial cells outside tumor regions as healthy epithelial:
Suppl. Fig 1: Inference of pretrained Hovernet, CellViT-256 and CellVit-SAM-H on TumSeg samples. We apply inference of state of the art models pretrained on Pannuke dataset, to 2 slides from TumSeg dataset for which tumor regions were annotated by 2 experts (marked in black). We display numbers of tumor and non-neoplastic epithelial cells detected by the different models. All inferences lead to prediction of tumor cells even outside tumor regions. The healthy epithelial cells (non-neoplastic epithelial cells) are poorly detected, and often scattered through the whole slide that they are not visible at the tissue level. Additionally, some visualization cannot be open (while no error during inference and enough memory provided, cell.geojson output from CellVit cannot open). We see that new strategies need to be implemented to discriminate better between tumor cells and healthy epithelial.
@@
where $I$ input Image containing $N$ pixels, $p_{i}(I;\mathbf{w_{0},w_{1}})$ regression output of HoVer branch at pixel $i$, $w_{0}$ and $w_{1}$ 2 sets of weights. $\Gamma_{i}$ defines the groundtruth of the horizontal and vertical distances of nuclear pixels to their corresponding centers of mass, the horizontal and vertical components of this map are denoted $\Gamma_{i,x}$ and $\Gamma_{i,y}$ respectively (see [14] for visualization and definitions). m denotes total number of nuclear pixels within the image and M denotes the set containing all nuclear pixels. $\nabla_{x}$ and $\nabla_{y}$ denote the gradient in the horizontal x and vertical y directions respectively. Finally, $X$ denotes the branch groundtruth, $Y$ the branch prediction, K is the number of classes and $\epsilon$ is a smoothness constant that was set to $10^{-3}$.
@@
On Table.1 the different hyperparameters tested are displayed. catmaxratiocat-max-ratio corresponds to the max area ratio that could be occupied by single category for a given crop of the input image. imgscaleimg-scale corresponds to the resizing of the image before cropping. These resizing can randomly be modified by a factor contained in [0.5, 2], following the data augmentation pipeline from the semantic segmentation library MMSegmentation [22].
Not all combinations were fully tested, some were aborted if the training loss did not decrease in the firsts epochs or if several hyperparameters in the set were already poorly performing in previous trainings. .
@@
Suppl. Fig 2: Example of distance overestimation with search area being a square. Here we have 1 cell in the squared search area $S$ and 2 cells in $(C\setminus S)$. While calculating the minimum distance of cells from the center $O$, if we only consider cells in $S$, we will take cellAcellA as being the closest cell. In reality the closest cell is cellBcellB. The distance calculated will be the distance between cellAcellA and the center, and will then be higher than the real closest distance. Nevertheless, the overestimation is bounded as expressed in Eq. 22 with $a=1$ as we are here in the case of a squared search area.
@@
Suppl. Fig 3: Probability of having distance overestimation representations, with search area being a square. (a) During the first search, if the search box is a square of side lenght 2r, we could have a distance calculation error if a cell is in the blue square $S_{1}$ without being in the pink circle $C_{1}$ (then in $S_{1}\setminus C_{1}$), and another cell is in black circle $C_{2}$ without being in in the blue square $S_{1}$ (then in $C_{2}\setminus S_{1}$). Then the cell in $C_{2}\setminus S_{1}$ could be missed from closest distance calculation and the cell in $S_{1}\setminus C_{1}$ considered as the closest cell. In other cases, such as having at least one cell in $S_{1}$, no distance overestimation could be made. (b) To be more precise, any cell in the hatched blue area inside $S_{1}\setminus C_{1}$ monitored by $x$ would lead for sure to an error if another cell is in the hatched black area inside $C_{2}\setminus S_{1}$ monitored by $x$ . For now we focused on the probability of overestimation in the case that at least one cell is found during the first search. If no cell is found, then the calculation differs for the further searching steps as visualized in (c) & (d). In such cases, a specific new area cannot contain any cells, as none were found during prior searches, impacting the full calculation.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\operatorname{area}(hatchblack)$ is the area of 4 circular segments of arc radius $d$ and sagitta ($d$ - $r_{1}$) then:
@@
where $P(A_{N})$ is calculated in (11), $P\bigl{(}U_{N}(0)\bigr{)}$ is calculated in (10), $P\bigl{(}U_{N}(n)\bigr{)}$ is calculated in (9) and finally $P\bigl{(}\bar{B_{N}}\mid U_{N}(n)\bigr{)}$ is developed in (13), (14), and (20).
@@
where $w$ is the width of the rectangle and $a$ the aspect ratio.
We covered here the case where at least one cell is found in the first search, meaning at least one cell is in $S_{1}$. In practice, distances can be overestimated in subsequent search iterations following the first. Unfortunately the calculation will be slightly different. Visualization of the cases where no cell is found inside $S_{1}$ are available at Supplementary Fig. 3(c), 3(d).
@@
We provide visualization of ROC curve for each splits.